William Pellico, Fermilab PMG, Nov 27 2013 3 FY13 Q4 Task Summary – Highlights
PIP FY13 Milestone Summary William Pellico, Fermilab PMG, Nov 27 2013 4 The current RLS is out of date and is being updated. Effort to adjust RLS when changes to budget and labor occur is challenging. Not Completed Completed FY13 Q1 FY13 Q2 FY13 Q3 FY13 Q4
WBS 1.1 LINAC FY13 Q4 Summary Tasks at a Glance 5 William Pellico, Fermilab PMG, Nov 27 2013 Linac WBS TaskFY13 Q4 Progress Summary WBS 22.214.171.124 High Level RF MBK design has been abandoned and substituted by a SBK specification. New bid process complete. Finalizing FP contract negotiation (warranty, T’s & C’s, etc…) WBS 126.96.36.199 Linac Modulator EE support completed 3 cell testing system. Implemented modifications to the design and initiated building 9 cell modulator WBS 188.8.131.52 7835 Procurement Waiting tube delivery procured in FY13 WBS 184.108.40.206 Simulations and Studies Small progress on this task during this quarter WBS 220.127.116.11 Linac Notch Creation Purchased of waveform generator and laser technology was chosen (1064nm Nd:YAG amplifier) WBS 18.104.22.168 Power Distribution Awaiting next long shutdown for L1 transformer installation WBS 22.214.171.124 Vacuum System Some equipment procured with Supplemental funds have been already delivered. Ongoing installation plans
6 William Pellico, Fermilab PMG, Nov 27 2013 WBS 1.1 LINAC (Highlights) WBS 126.96.36.199 Linac Notch Creation Linac Notch – Beam Shaping Laser technology chosen 1064nm technology based upon Nd:YAG amp offered by Northrop Grumman Cutting Edge Optical (NGCEO) Started with Notcher installation plans and stand design Old New
3-cell simultaneous turn-on at 300 amps. WBS 188.8.131.52 Modulators No problems with 3-cell operation as long as initial turn on timing is done properly. Simultaneous turn-on yields 2700 volts with 250 ns rise time (10% - 90%). 9-cell construction under-way William Pellico, Fermilab PMG, Nov 27 2013 7
8 WBS 1.1 LINAC WBS 184.108.40.206 High Level RF Klystron Timetable MBK 2012 Evaluated LE Linac RF power situation (wrote document - recommending Klystron option) Wrote specifications Bid process (1 bid – CPI) End of year funding issue Contract (CPFF) MBK 2013 Carryover funding issue Late 2013 – carryover approved Contract type causes concern Lots of negotiations Re-issue bid for SBK SBK 2014 New bid received Fix Cost Negotiating contract issues (CPI)
WBSTaskHighlight 220.127.116.11 Anode Supply Supplemental funding allowed initial engineering to start with purchase of internal transformer 18.104.22.168 Cavity Refurbishment 8 th & 9 th cavities refurbished & certified (Aug 30 & Oct 21); however, 1 st refurbished cavity removed due to RF leak – replaced a tuner and refurbished flange connections 22.214.171.124 New Tuners Delivery of nearly 200 low mu cores – pass Await delivery latest round of high mu cores – hope is that vendor will have increased mu 126.96.36.199 Spare RF Cavity Old rejected pieces reworked with some new components to make a vacuum certified cavity; requires 3 tuners 188.8.131.52 Simulations & Studies Team working on optics software process/procedure 184.108.40.206 Notcher System Operational – working on short kickers/PS 220.127.116.11 Power Distribution Supplemental funding moved ordering new transformer earlier 18.104.22.168 Vacuum Systems Supplemental funding moved ordering replacement vacuum equipment earlier William Pellico, Fermilab PMG, Nov 27 2013 9 WBS 1.2 Booster FY13 Q4 Summary Tasks at a Glance
William Pellico, Fermilab PMG, Nov 27 2013 10 WBS 1.2 Booster (Highlights) 22.214.171.124 Notcher System 126.96.36.199 Cavity Refurbishment (more later) Notcher System Testing of Short Kickers: Magnets passed impedance testing Supply nearing completion Repair of 9 th cavity Repair LCW cone leakRefinished outer flange Center tube, fingers (not shown) and stem burnt New short kicker pulser Two short kickers
Low mu ferrites cores for new tuners 185 cores from National Magnetics has been received in three different shipments; in Sept, Oct and Nov. The following data was taken for all cores including a low power ƒo measurement. Core # muOD (cm)ID (cm) Thickness cmgrams volume cm^3 density gm/cm^3Qofo MHzf lowf high M4-2013-10011.17 20.3212.702.542111503.024.20274.143.58843.50943.668 M4-2013-10111.76 20.3112.702.542179502.254.34285.343.07943.00443.155 M4-2013-10211.55 20.3112.702.542158501.784.30292.343.26543.19243.340 M4-2013-10311.65 20.3112.702.542172502.044.33282.243.17543.09943.252 William Pellico, Fermilab PMG, Nov 27 2013 11
High Power Test 30 cores from across all lot numbers were tested in the high power test jig with 10,000 Amps of DC bias and driven with a 5-KW RF amplifier. Had very consistent results. No high loss effect (HLE) in any of the cores. M4-2013-107Lot 56371-1 Bias Current (A) Setting (dBm) Res. Freq. (MHz) Forward Power (V) Reverse Power (V) Forward Power (kW) Reverse Power (W) Cavity Voltage 07.041.0121.2000.2603.5991696.330 30007.042.3301.2000.3003.5992259.200 50007.042.6071.2200.2703.72018210.500 100007.042.9101.1800.3503.48030611.150 Q measurement Bias Current (A)300010000 SG setting (dBm)7.0 Max Cavity Voltage (V)9.211.15 Resonant frequency (MHz)42.33042.910 f142.29142.879 f242.37642.940 Q 498.0703.4 William Pellico, Fermilab PMG, Nov 27 2013 12
William Pellico, Fermilab PMG, Nov 27 2013 17 PIP Labor Usage (FTE) Planned Actual
William Pellico, Fermilab PMG, Nov 27 2013 18 Planned Actual
Supplemental Spending – Direct M&S 19 ID PIP Tasks Task Code 1 DTI switches and controller for extractor (and spare for Einzel Chopper) 2103.01.01.02.02.AD 2 Booster Transformer (MCC and spare breaker) 2103.02.05.02.01.AD 3 Buncher Solid State RF power 2103.01.01.02.02.AD 4 12 Scroll Pumps for Booster 2103.02.06.01.AD 5 Linac Roughing Station 2103.01.05.03.AD 6 Laser Amp for Linac Notcher 2103.01.02.03.AD 7 1 st Linac 7835 188.8.131.52.1 8 RFQ Turbo and Roughing station/controls 2103.01.06.01.AD 9 400 MeV line pumps 2103.01.05.03.AD 10 2 nd Linac 7835 184.108.40.206.1 11 14 Booster roughing and turbo pumps stations 2103.02.06.01.AD 12 Replace 5 Booster valves 2103.02.06.01.AD 13 Linac Modulator (SLAC Version) 2103.01.01.02.01.AD 14 SS amp for Linac Driver system 2103.01.01.02.02.AD 15 Booster Anode Supply Transformers 2103.02.01.01.02.AD 16 Laser notch pulser 2103.01.02.03.AD 17 201 MHz Circulator 2103.01.01.01.02.AD 18 Linac 55 chilled water pump and plumbing 2103.01.05.02.AD 19 3 rd Linac 7835 2103.01.01.04.AD 20 Injector Solenoid and trim PS 220.127.116.11.8 21 RF test ‘cave’ or shielded room 2103.02.01.05.02.AD Supplemental funds were provided in July – This is a significant $$ and will have a positive effect on the schedule. Part of funds will be used to remedy the recent PS budgets which impacted PIP goals (7835 tubes). Part of funds will be used to put back track PIP items impacted from previous reductions (e.g. laser notcher). Part of funds will be used to move tasks forward (Refurb) Some small items already delivered – all should be in FY15 William Pellico, Fermilab PMG, Nov 27 2013
20 PIP M&S Notes/Comments This includes carryover and the supplemental funding received in August.
William Pellico, Fermilab PMG, Nov 27 2013 23 Operations with refurbishment means the voltage and loses are always a concern – medium intensities. Batch intensity will track the voltage sum – mostly due to transition loses. Same scales for 2011 operations and present operations. Average voltage is about 817 KV compared to 917 two years ago. This is likely to continue with occasional unplanned downtime to fix RF stations.
PIP FY13 Q4 Summary & Concerns WBS 1.1 Linac Linac WBS efforts are in good shape Klystron effort continues – Business Services/CPI/PIP Switching from MBK to SBK Modulator efforts continue with both SLAC and AD/EE support Simulation and Studies-> continue development code for DTL lattice, WBS 1.2 Booster Cavity Refurbishment – Labor & Time constraints Contract Employees Dosage (ALARA) – will need to rotate personnel Need to deal with operations vs. repair (MI startup) Booster shielding – progressing Booster has not found a ferrite that meets our High μ need A order has been placed for 10 high u ferrites – getting close to a usable ferrite FY14? Booster High Intensity/rep rate: Up to 7 Hz – MiniBooNE, Recycler and NoVA Booster Rep rate limit – RF has had more downtime (rep rate or operating time?) Supplemental Funding – Big Boost and will help PIP/FY14 William Pellico, Fermilab PMG, Nov 27 2013 24